OpenStudy (anonymous):

Write out the steps to solve the equation 3n = 64 - 5n

OpenStudy (anonymous):

1st we want to get all of the n's on one side. To do that add 5n to both sides of the equation. Now what do we have?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

then

OpenStudy (anonymous):

What equation to you have now?

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

5=69-n

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

No....add 5n to both sides, like this 3n = 64 - 5n +5n +5n 8n = 64 <----This is the equation we have after step 1. Now, what do we do to both sides to get rid of the 8?

OpenStudy (anonymous):

i dont no thats why im asking u

OpenStudy (anonymous):

- 8 from oth sides

OpenStudy (anonymous):

No, not -8. Divide both sides by 8. That will isolate the n on the left side and we shall have the answer. What do you get when you divide both sides by 8?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

1n = 8

OpenStudy (anonymous):

n = 8 You're finished

OpenStudy (anonymous):

thnx